OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x3 + x2 - 42x

OpenStudy (anonymous):

0 would be one, because if you plug 0 in for x, it would all equal 0

OpenStudy (anonymous):

so factor x out and then try finding the factors of the function f(x)=x^2+x-42

OpenStudy (***[isuru]***):

factorize it... x^3 + x^2 -42x = x ( x^2 + x - 42 ) = x ( x^2 + 7x - 6x - 42 ) = x [ x ( x + 7) - 6( x + 7) ] = x [ (x - 6) ( x + 7) ]

OpenStudy (lieutenantgeneral):

f(x) = x(x+7)(x-6) zeros are: -7, 0, 6

OpenStudy (***[isuru]***):

now equal it to zero and solve it for x... and u vill have ur zeros x [ ( x -6) ( x + 7) ] = 0 which means x = 0 or x = 6 or x = -7 so.. the zeros r -7 , 0 , 6